OpenStudy (anonymous):

Simplify -1 3/7 * (-3 2/3)

OpenStudy (anonymous):

-1*(3/7)=-3/7 -3*(2/3)=-2 -3/7*-2=6/7

OpenStudy (anonymous):

Can you explain how to solve this equation please ?